OpenStudy (sissyedgar):

Gary adds $12 to his savings each week. At Week 3, Gary had $60 total in his savings. How much will Gary have saved after 15 wk? $204 Right?

OpenStudy (sissyedgar):

@misty1212

OpenStudy (malcolmmcswain):

204 is correct @sissyedgar

OpenStudy (sissyedgar):

thx!

OpenStudy (malcolmmcswain):

12*12=144+60=204